OpenStudy (cedric):

Choose the Standard Form of the equation y = -1/3x - 4 1/3x + y = 4 3x + y = −4 x + 3y = −12 −3x + y = −12

OpenStudy (anonymous):

standard form is ax+by=c

OpenStudy (cedric):

oh so it has to be A thnak you!

OpenStudy (anonymous):

Yessir:)

OpenStudy (cedric):

i see isee so would i do that for every problem like this?

OpenStudy (anonymous):

AH! My bad, you would multiply the equations out. y = -1/3x - 4 then add 1/3 to both sides so 1/3x+y=-4 then multiply by 3 x+3y=-12

OpenStudy (anonymous):

x can't be a fraction!

OpenStudy (cedric):

Why cant it be a fraction though? is it because of the denomenator??

OpenStudy (anonymous):

Because standard form can only contain integers:P

OpenStudy (cedric):

ohhhhh